The perimeter of a rectangle is 56 inches. the ratio of the length to the width is 6:1. find the length and the width
how would i do this?

Found 2 solutions by mananth, rapaljer:
Answer by mananth(16946)

The perimeter of a rectangle is 56 inches. the ratio of the length to the width is 6:1. find the length and the width
Ratio is 6/1
multiply by common factor x
the length = 6x & width = 1*x
perimeter = 2(L+W)
Perimeter = 2(6x+x)
56=14x
/14
x=4
width = 4
and the length = 24
CHECK
24+24+4+4 = 56

Answer by rapaljer(4671)

Let x=width
6x=length

2W+2L=P
2(x) + 2(6x)=56
2x + 12x = 56
14x = 56
x=56/14
x= 4 inches Width
6x=24 inches Length

Check: 2W + 2L = P
2(4) + 2(24) = 56
8 +48 =56 It checks!!